a rock is sliding down a hill. Assume that the hill is 20 meters high and the rock hs a mass of 1000 kilograms. If the rock started at the top of the hill with no initial speed, what is its speed when it reaches the bottom of the hill? use Gn 10m/s^v Give answer in kilojoules.
Answer #1

By energy conservation,

KEi + PEi = KEf + PEi

here, KEi = 0 (because no initial speed)

PEi = m*g*hi

KEf = 0.5*m*v^2

PEf = m*g*hf

given, g = 10 m/s^2

hi = 20 m

hf = 0

So, 0 + m*g*hi = 0.5*m*v^2 + m*g*hf

v = sqrt(2*10*20)

v = 20 m/s

Please upvote.